Ordered: Lidocaine 4 mg/minute

Available Infusion pump. Standard solution of 4 g in 500 ml D5W

What is the pump setting?
Answer #1

ANSWER.

* Calculation of infusion rate(ml/hr).

          ml/hr = Volume(ml) / Time(hr).

    Calculation of volume.

         Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

       Desired dose(D) = 4 mg/min.

       Dose in hand(H) = 4000 mg (4g).

           Quantity(Q) = 500 ml.

As per above formula

      Volume(ml) of drug required = D/H * Q

                                                    = 4/4000 * 500

                                                    = 0.001 * 500

                                                    = 0.5 ml.

                             Time = 1 min.

Convert ml/min to ml/hr by multiply with 60

                     ml/hr = 0.5 * 60

                              = 30 ml/hr.